1.)Which of the expressions is equivalent to the following statement:

The sum of two even numbers is even.

a.) If x is even or y is even, then x + y
b.) If x is even or y is even, then x + y is even
c.) If x is even and y is even, then x + y is not even.
d.) If x is even and y is even, then x + y is even

2.) Find a counterexample for the statement: every month of the year has 30 or 31 days.

a.) March
b.) December
c.) January
d.) February

3.) If direct proof is used to prove the following statement: If x is a real number and x ≤ 3, then 12 - 7x + x*x ≥ 0. What is the hypothesis?

a.) If x is not a real number or x > 3
b.) If x is a real number and x ≤ 3
c.) 12 - 7x + x*x < 0
d.) 12 - 7x + x*x ≥ 0

4.) If proof by contrapositive is used to prove the following statement: For every real number x, if x is irrational, then -x is also irrational. What is the hypothesis?

a.) x is rational.
b.) x is irrational.
c.) -x is rational.
d.) -x is irrational.

5.) Consider the following statement: For every positive integer n less than 3, n + n ≥ n2. Which facts must be checked in a proof by exhaustion of the statement?

a.) n = 0, 0 + 0 = 0; n = 1, 1 + 1 > 1
b.) n = 1, 1 + 1 > 1; n = 2, 2 + 2 = 22
c.) n = 0, 0 + 0 = 0; n = 1, 1 + 1 > 1; n = 2, 2 + 2 = 22
Answer #1

1(d) If x is even and y is even, then x + y is even

Explanation: BOTH should be even for the sum to be even

2(d) February
Explanation: As we know it can have 28/29 days


3)b.  If x is a real number and x ≤ 3

4)a. x is rational.

5) c. n = 0, 0 + 0 = 0; n = 1, 1 + 1 > 1; n = 2, 2 + 2 = 22


Thanks, let me know if there is any concern